1. Sketch the region between ?=?3+1, y=0, x=1 and x=2. Find the volume of the solid obtained when you rotate it around the y-axis.
Answer #1

Solution: Given curves are y = x + 1, y = 0,X = 1 and x = 2 7=X Shells Method (about the y-axis) Volume, V = 21 * xf(x)dx Her